The number of flaws per square yard in a type of carpet material varies with mean 1.4 flaws per square yard and standard deviation 1 flaws per square yard. This population distribution cannot be normal, because a count takes only whole-number values. An inspector studies 171 square yards of the material, records the number of flaws found in each square yard, and calculates x, the mean number of flaws per square yard inspected. Use the central limit theorem to find the approximate probability that the mean number of flaws exceeds 1.5 per square yard. (Round your answer to four decimal places.)
Answer #1

By Central Limit Theorem, even though population distribution is not Normal Distribution, the sampling distribution of sample mean is Normal Distribution with mean = mean of the population = \mu = 1.4 and Standard Deviation = SE = \sigma /\sqrt{n} = 1/\sqrt{171} = 0.0765.

To find P(\bar{x} > 1.5):

Z = (1.5 - 1.4)/0.0765 = 1.3077

Table of Area Under Standard Normal Curve gives area = 0.4049

So,

P(\bar{x}>1.5) = 0.5 - 0.4049 = 0.0951

So,

Answer is:

0.0951